### v1

@@ Abstract
Optimizing molecules to improve their properties is a fundamental challenge in drug design. For a fine-tuning of molecular properties without losing bio-activity validated in advance, the concept of bioisosterism has emerged. Many in silico methods have been proposed for discovering bioisosteres, but they require expert knowledge for their applications or are restricted to known databases. Here, we introduce DeepBioisostere, a deep generative model to design suitable bioisosteric replacements. Our model allows an end-to-end chemical replacement by intelligently selecting fragments for removal and insertion along with their attachment orientation. Through various scenarios of multiple property control, we showcase the model’s capability to modulate specific properties, addressing the challenge in molecular optimization. Our model’s innovation lies in its capacity to design a bioisosteric replacement reflecting the compatibility with the surroundings of the modification site, facilitating the control of sophisticated properties like drug-likeness. DeepBioisostere can also provide previously unseen bioisosteric replacements, highlighting its capability for exploring diverse chemical modifications rather than just mining them from known databases. Lastly, we employed DeepBioisostere to improve the sensitivity of a known SARS-CoV-2 main protease inhibitor to the E166V mutant that exhibits drug resistance to the inhibitor, demonstrating its potential application in lead optimization.
@@ 1 Introduction
Optimizing molecules to fulfill multiple properties is a challenging yet fundamental goal of drug design. A drug should be bioactive, bio-compatible, and synthetically accessible,[1] taking an enormous time and cost to bring a compound from a lab to a patient.[2] One promising approach to achieving this multi-objective molecular design problem is modifying a chemical structure from an acclaimed molecule with satisfied properties. Namely, chemical modifications, such as scaffold hopping[3, 4] or R-group replacement,[5] aim to delicately enhance a target property while maintaining the other properties of the original structure, where a comprehensive consideration of both structural and biochemical analogies is necessary. To develop a more efficient way of chemical modification, many chemists attempted to standardize useful rules that are frequently adopted in molecular optimization, leading to the emergence of the concept of bioisosteric replacement.[6, 7, 8] A number of chemical moiety pairs involved in bioisosteric replacements, so-called bioisosteres, have been reported for the last decades and facilitated more proper chemical modifications.[9, 10, 11]
@@
Traditional bioisosteric replacements heavily relied on the intuition and experience of medicinal chemistry expertise.[12] Identification of proper bioisosteric replacements has required extensive trial and error, leading to the rise of in silico approaches.[13, 14] There have been two mainstream in silico approaches to finding bioisosteres—similarity-based and database mining methods. Similarity-based approaches operate on the assumption that structural analogs would exhibit similar biological or chemical properties.[15] They prioritize potential bioisosteres by comparing their structural, steric, or electronic features to the original moiety of interest.[16, 17, 18, 19] Diverse molecular descriptors such as fingerprints[20, 21] or structural alignments[22] were often employed to assess the similarity between chemical moieties. The similarity-based methods facilitated the expectation of biochemical property changes during a particular chemical modification before proceeding to experimental validation.
@@
On the other hand, database mining approaches identify potential bioisosteres through the extraction from large chemical databases.[23, 24, 25] (see Fig. 1) They systematically enumerate pairs of molecules with only a small structural difference, the so-called matched molecular pairs(MMPs),[26] from a public biochemical library such as ChEMBL.[27] Then, molecular fragments corresponding to the difference in each MMP are paired and identified as potential bioisosteric replacements along with their occurrences in the parent library. The differences in chemical properties and bioactivity in biochemical assays are also used to expect the effect of each bioisosteric replacement, which is essential for subsequent bioisostere recommendations. This statistical evidence enables the identification of a new bioisosteric replacement, even if the involved moieties are structurally unrelated at first glimpse.
Despite each approach’s advantages, limitations still exist and hamper the identification of a proper chemical modification. The similarity-based methods highly depend on the correlation between carefully crafted molecular descriptors and a specific property to be controlled. Yet, finding suitable descriptors from a variety of available options[28] inevitably relies on expertise and is often arbitrary. The similarity-based methods may also struggle to identify bioisosteres that are structurally dissimilar but able to modulate particular molecular properties in the desired way. The database mining methods, on the other hand, cannot assess bioisosterism between unseen chemical moiety pairs that do not appear in the parent database due to their statistical absence. Also, the expectation of property changes based on the mined data may not work for more complicated properties that depend on complex relationships between internal moieties, such as synthesizability or drug-likeness. Moreover, although both methods propose potential bioisosteres once the target molecular fragment is provided, the decision on which substructure to modify still relies on the experts’ assessment of target properties and molecular structures. Since the design of bioisosteric replacements involves selecting which molecular fragment to modify and determining its bioisosteres, neither the similarity-based nor database mining approaches can fully address the molecular property control process based on bioisosteric replacements.
Recently, deep learning-based chemical modification approaches have stood out to featurize data-intrinsic information that contains chemical and biological contexts without handcrafted descriptors.[29, 30, 31, 32, 33, 34, 35] Ertl first used a deep learning approach in identifying potential bioisosteres.[29] In that work, he constructed training data from ChEMBL and utilized substructural information to train a neural network on the inherent relationship between analogous fragments. Remarkably, this model could propose bioisosteric relationships not present in the training dataset, serving as novel bioisostere candidates. In parallel, a number of approaches that directly model the chemical modification task with deep generative models have appeared. They regard the chemical modification problem as a language translation[30, 31, 32] or a graph-editing task[33, 34, 35] by representing a molecule as a text or a molecular graph, respectively. The advantage of applying deep generative modeling is that multiple property changes can be accounted for designing chemical modifications, i.e., changing solubility while maintaining molecular similarity. Furthermore, they could identify novel modified structures with improved biochemical properties by learning the intrinsic relationships between molecular structures and their analogs. Based on these advantages, deep generative modeling is expected to significantly contribute to overcoming the limitations of the conventional approaches for identifying and prioritizing bioisosteres.
Here, we propose DeepBioisostere, a deep generative model for multi-property control with designing suitable bioisosteric replacements in an end-to-end manner. Our model optimizes a molecular structure in a two-step strategy: selecting a substructure to be removed and inserting a proper alternative moiety, i.e. bioisostere, to achieve target properties. With the sequential modeling of bioisosteric replacement, our model designs the overall process of chemical modification while the previous similarity and database mining approaches aim to provide bioisosteres for a moiety selected by expertise in advance. Our step-wise strategy enables more flexible suggestions of novel bioisosteric replacements even if they are not included in training data. By utilizing a deep learning-based approach for chemical modification, DeepBioisostere could control complicated properties such as synthetic accessibility and drug-likeness simultaneously despite the difficulty due to their dependency on an overall molecular structure. Also, we demonstrate that the compatibility between an insertion moiety and its surroundings to be attached is crucial for multiple property control, which can be comprehensively considered by DeepBioisostere. Lastly, we adopted DeepBioisostere to the challenging problem of treating drug resistance due to a point mutation of SARS-CoV-2 main protease, implying its potential application in the lead optimization process. To our best knowledge, this is the first work that proposes an in silico method to design the complete procedure of bioisosteric replacement without any aid of expertise, facilitating an autonomous lead optimization process. (see Fig. 1)
@@ 2.1 Brief introduction of DeepBioisostere
The DeepBioisostere model is trained on an MMP dataset with structural analogs with small differences, which we constructed from the ChEMBL database. Since we designed the model with three main modules corresponding to each step for the chemical modification, it learns proper ways of removing and then inserting fragments to meet the given condition of multiple biochemical properties. This sequential modification facilitates the optimization of a molecule while retaining the overall properties except the target ones to be improved. Notably, the DeepBioisostere model comprises two embedding networks at different levels—atom and fragment. This unique feature, especially the latter network, allows the model to learn the compatibility between an insertion fragment and the remaining parts, enabling the fine control of complex properties depending on a whole molecular structure. For synthetically accessible chemical modification, the DeepBioisostere model enumerates molecular fragments inside a given molecule based on the breaking retrosynthetically interesting chemical substructures(BRICS)[36] rules and selects one of them as a removal fragment. Then, the model selects an insertion fragment from a pre-defined fragment library by examining all of its items in terms of their compatibility with the remaining part of a given molecule and the target property control condition. More details about the training and generation processes of the DeepBioisostere model can be found in Section 4.
@@ 2.2 Multi-property Control with DeepBioisostere
We first apply our DeepBioisostere model to multi-property control tasks. During the optimization of a potent molecule in drug development, its structure is modified and elaborated to improve biochemical properties, for example, lipophilicity, water solubility, synthetic accessibility, and drug-likeness. This so-called multi-property control is more challenging than controlling a single property since, for example, the lipophilicity can be increased simply by elongating a carbon chain, but the modified structure is often not drug-like.[37] Thus, a comprehensive consideration of the whole ligand structure and its properties is required to modulate multiple properties simultaneously.
@@
Here, the goal is to provide modified molecular structures to meet a given multi-property condition. We tested our model for three scenarios that practitioners might encounter in a lead optimization process: (1) increasing or decreasing ligand lipophilicity while maintaining its molecular size, (2) increasing drug-likeness while maintaining its molecular weight, and (3) alleviating the synthetic complexity of a ligand with fair bio-activities but low synthetic accessibility. Obviously, the second and third scenarios are more challenging than the first one, which requires more comprehensive considerations about the relationship between the properties of interest. Throughout Section 2.2, the following criteria are commonly employed for selecting test molecules: (1) a single neutral molecule, (2) absence of more than three rings, (3) inclusion of at least one BRICS bond, and (4) molecular weight falling within the range of 250 to 500. In addition, to examine our model’s capability to find novel bioisosterism, we only used fragments from the test dataset for insertion, which were never adopted as ground truths nor negative samples for the insertion step in the training process.
@@
The overall property control results are exhibited in Table 1. For scenario 1, we selected 100 unique molecules whose log are within the range of 1 to 5, regarded as moderate lipophilicity. Whether one should increase a molecule’s log or not depends on its specific application; for example, the biological target under consideration or the drug delivery plan are critical factors in making such a decision. Thus, we tested our model for both cases with two property control conditions: increasing log by 1 and decreasing log by 1 while maintaining overall molecule size in terms of molecular weight(MW). We note that log is a logarithm of the octanol-water partition coefficient, so the difference by 1 corresponds to a 10 times larger or smaller solubility ratio.[38] We generated 100 unique molecular structures for each original molecule according to the two property control conditions. For a few test molecules, DeepBioisostere generated less than 100 molecular structures for each one since the BRICS-allowed insertion fragments were less than 100. The results are summarized in the first row on Table 1. We noticed that the mean values of the molecular weight differences between an original molecule and its modified ones were +6.60 and +3.34, respectively. These values suggest that the molecular weights of most modified structures are different from the original ones by less than one heavy atom. (see Supplementary Figure 4 for more details of property distributions) Meanwhile, log was modulated with averaged changes of +0.81 and -0.86, falling slightly behind the requested property values. These statistics indicate that our model successfully adjusted the log values of the given molecular structures with barely changing their molecular weights, which requires careful selection of both removal and insertion fragments.
@@
Scenario 1 | 0 | +6.60 | 14.0 | |
log | +1 | +0.81 | 0.35 | |
0 | +3.34 | 13.0 | ||
log | -1 | -0.86 | 0.37 | |
Scenario 2 | 0 | +2.73 | 12.7 | |
QED | +0.1 | +0.064 | 0.098 | |
0 | +1.90 | 12.4 | ||
QED | +0.2 | +0.087 | 0.101 | |
Scenario 3 | QED | 0 | -0.049 | 0.092 |
SAscore | -0.5 | -0.294 | 0.242 | |
QED | 0 | -0.050 | 0.09 | |
SAscore | -1.0 | -0.501 | 0.224 |
Second, we demonstrated an optimization process where a molecular structure is modified to improve its drug-likeness in terms of QED.[39] (scenario 2 in Table 1) In this scenario, only molecules whose QED values lie in the range of 0.4 to 0.6 were used as test molecules. Property control conditions were set to increase the QED value by 0.1 or 0.2 while keeping molecular weight constant. By employing the DeepBioisostere model, we were able to obtain novel molecules with enhanced drug-likeness; the mean values of QED changes were +0.064 and +0.087, respectively. QED is calculated based on eight physicochemical descriptors, including the molecular weight. Thus, finely modulating QED while preserving molecular weight is challenging since the other descriptors, such as the number of hydrogen bond donors and acceptors, rotatable bonds, and aromatic rings, should be adjusted in the desired way. We note that the result in scenario 2 implies that our model can capture information to modulate those descriptors for increasing the QED value. Although the property control condition specifying a larger QED increase yielded molecules with higher QED values, deviation from the given condition also became larger. It is not surprising that increasing QED by 0.2 is more challenging since its occurrence within the training dataset is rarer. (see Supplementary Figure 2)
Lastly, we tested our model to improve the synthetic accessibility of a molecule that has a high QED value but low synthetic accessibility. (scenario 3 in Table 1) Here, we adopted the SAscore introduced by Ertl and Schuffenhauer [40], a well-known estimator of the synthetic accessibility. We collected test molecules for this scenario with criteria that QED value larger than 0.7 and SAscore larger than 4. Note that SAscore of a simple catalog molecule is about 2 or 3, whereas for a natural product, it is mostly greater than 5.[40] Property control conditions were set to decrease the SAscores of molecules by 0.5 or 1.0 while retaining their QED values. In this scenario, our model provided novel molecules with lower SAscores, although several of them exhibited slightly decreased QED values. This result suggests that DeepBioisostere can be adopted to supplement many in silico methods for drug design whose outcomes are often questioned in their synthetic accessibility.[41, 42]

### v2

@@ Abstract
Optimizing molecular properties while preserving biological activity is a central challenge in drug design. Bioisosteric replacement, which substitutes a molecular fragment with a chemically or biologically analogous moiety, offers a powerful strategy for fine-tuning properties without disrupting target binding. However, existing in silico approaches often rely on expert-defined modification sites or suffer from modulating multiple molecular properties simultaneously. Here, we present DeepBioisostere, a deep generative model that performs end-to-end bioisosteric replacement by autonomously selecting and substituting molecular fragments to satisfy multiple target properties. The model captures complex relationships across the molecular graph, enabling the optimization of sophisticated properties such as drug-likeness and synthetic accessibility. By learning from experimental bio-assay data, DeepBioisostere proposes replacements that maintain biological activities, even generating potential bioisosteres beyond the training data. We demonstrate the effectiveness of the model in computational hit-to-lead optimization scenarios, highlighting its potential to accelerate rational molecular design without relying on expert heuristics or pre-established substitution rules.
@@ 1 Introduction
Optimizing molecules to fulfill multiple properties such as bioactivity, bio-compatibility, and synthetic accessibility simultaneously is a fundamental yet challenging goal in drug design, often requiring substantial time and cost to advance a compound from the laboratory to the clinic.[drugproperty] Given the complexity of the multi-objective design problem, a widely adopted approach is to begin with a molecule that exhibits known bioactivity toward a target protein and gradually modify its structure to enhance desirable properties. In particular, bioisosteric replacement, the substitution of one chemical moiety with another that preserves biological activity, offers a principled way to fine-tune molecular properties with minimal disruption to the molecule’s overall function.[Bioisostere1, Bioisostere2, Bioisostere3] This approach allows for delicate enhancement of specific molecular properties while maintaining the essential characteristics of the original compound, requiring careful consideration of both structural and biochemical analogies. To systematize this process, many chemists have attempted to formalize common patterns observed in successful molecular replacement, namely bioisosteres. In recent decades, numerous bioisosteres have been reported, facilitating more rational and effective chemical modifications in drug development.[example_tetrazole1, example_tetrazole2, phenyl-bioisosteres]
@@
Traditional bioisosteric replacements have heavily relied on the intuition and experience of medicinal chemists.[patani1996bioisosterism, ertl2007silico, kumari2020amide, seddon2018bioisosteric] The identification of proper bioisosteric replacements has required extensive trial and error, leading to the rise of in silico approaches.[InSilicoApproach, SwissBioisostere1] There have been two mainstream in silico approaches to finding bioisosteres: similarity-based and database-mining methods. Similarity-based approaches operate on the assumption that structural analogs would exhibit similar biological or chemical properties.[SimilarityPropertyPrinciple] They prioritize potential bioisosteres by comparing their structural, steric, or electronic features to those of the original moiety of interest.[SimilarityApproach1, SimilarityApproach2, SimilarityApproach3, SimilarityApproach4] Diverse molecular descriptors such as fingerprints[example_similarity, example_similarity2] or structural alignments[example_structure] were often employed to assess the similarity between chemical moieties. The similarity-based methods facilitated changes in biochemical properties during a particular chemical modification prior to experimental validation.
@@
On the other hand, database-mining approaches identify potential bioisosteres by extraction from large chemical databases.[DatabaseMiningApproach1, DatabaseMiningApproach2, DatabaseMiningApproach3] (See Fig. 1) They systematically enumerate pairs of molecules with only a minor structural difference, the so-called matched molecular pairs (MMPs),[MMP] from a large bio-assay database provided in public libraries such as ChEMBL.[ChEMBL] Then, molecular fragments corresponding to the difference in each MMP are paired and identified as potential bioisosteric replacements along with their occurrences in the parent library. The differences in chemical properties and bioactivity in biochemical assays are used to predict the effect of each bioisosteric replacement, referred to as matched molecular pairs analysis (MMPA), which is essential for subsequent bioisostere recommendations. This statistical evidence enables the identification of a new bioisosteric replacement, even if the involved moieties are structurally unrelated at first glance.
Despite their advantages, existing approaches face limitations that hinder the identification of suitable chemical modifications. Similarity-based methods rely heavily on predefined molecular descriptors correlated with specific properties, making them dependent on expert intuition and prone to arbitrariness.[NumDescriptors] They also struggle to identify bioisosteres that are structurally dissimilar but functionally equivalent. Meanwhile, database-mining methods are constrained by the statistical absence of uncommon or unexplored moiety pairs within the database and thus can barely identify them. Because such methods evaluate property changes based only on the exchanged fragments, they struggle to capture complex attributes such as synthesizability or drug-likeness, which inherently depend on the entire molecular context rather than local substitutions. Moreover, both approaches require the manual selection of the molecular fragment to be modified, which limits fully automated molecular optimization through bioisosteric replacement.
Recently, deep generative modeling has emerged as a powerful approach for chemical modification, capturing data-intrinsic patterns that reflect both chemical structure and biological context without relying on handcrafted descriptors.[FindBioWithDNN, shan2020molopt, jin2020multi, fu2021mimosa, loeffler2024reinvent, TransformerMMP1, TransformerMMP2, TransformerADMET, jin2019hierarchical, jin2020hierarchical, Modof, DeepHop, ScaffoldGVAE, maziarka2020mol] For example, FindBioWithDNN and shan2020molopt applied neural networks to fragment-level data derived from ChEMBL, demonstrating the ability to identify potential bioisosteric replacements beyond the training set. jin2020multi, fu2021mimosa, and loeffler2024reinvent employed standard optimization algorithms to refine molecular structures toward specific biological targets with low-cost labels, such as docking scores. DeepHop and ScaffoldGVAE demonstrated that deep learning methods can be effectively applied to scaffold hopping, generating alternative scaffolds while maintaining biological activity. While the methods mark significant progress, no existing deep learning framework for chemical modification can simultaneously control multiple properties while maintaining the original bioactivity in a general manner. Addressing this gap is essential for fully leveraging bioisosteric replacement in automated, property-driven molecular design.
Here, we propose DeepBioisostere, a deep generative model for autonomous bioisosteric replacement with fine control over multiple molecular properties in an end-to-end manner, selecting modification sites and generating context-aware substitution. (See Fig. 1) We systematically curated a standardized dataset for training with experimental bioassay data to implicitly incorporate the statistical patterns of structure-activity relationships, guiding our model’s proposed replacements toward maintaining the target activity. DeepBioisostere can address the compatibility of replacement with the whole molecular structure, which is necessary to optimize delicate molecular properties such as synthetic accessibility and drug-likeness. Furthermore, we demonstrate the ability of our model to identify unexplored bioisosteric replacements that achieve the target property, which is beyond the capacity of conventional approaches. Finally, we apply DeepBioisostere in an in silico hit-to-lead workflow and demonstrate its ability to improve key biochemical properties of hit candidate molecules while maintaining binding affinity, offering a new opportunity toward automated drug design.
@@ 2.1 Brief Introduction of DeepBioisostere
The DeepBioisostere model is trained on an MMP dataset with structural analogs with minor differences, which we constructed from the ChEMBL database. Since we designed the model with three main modules corresponding to each step in the chemical modification process, it learns the proper ways to remove and then insert fragments to meet the given conditions of multiple biochemical properties. This sequential modification facilitates the optimization of a molecule while retaining its overall properties, except for the target ones to be improved. Notably, the DeepBioisostere model comprises two embedding networks at different levels: atom and fragment. This unique feature, especially the latter network, enables the model to learn the compatibility between an insertion fragment and the remaining parts, allowing for fine control of complex properties that depend on the entire molecular structure. For synthetically accessible chemical modification, the DeepBioisostere model enumerates molecular fragments within a given molecule based on Breaking Retrosynthetically Interesting Chemical Substructures (BRICS)[BRICS] rules, and selects one of them as a removal fragment. Then, the model selects an insertion fragment from a pre-defined fragment library by examining all of its items in terms of their compatibility with the remaining part of a given molecule and the target property control condition. More details about the training and generation processes of the DeepBioisostere model can be found in Section 4.
@@ 2.2 Multi-property Control with DeepBioisostere
We first apply our DeepBioisostere model to multi-property control tasks. During the optimization of a potent molecule in drug development, its structure is modified and elaborated to improve biochemical properties, such as lipophilicity, water solubility, synthetic accessibility, and drug-likeness. This so-called multi-property control is more challenging than controlling a single property since, for example, the lipophilicity can be increased simply by elongating a carbon chain, but the modified structure is often not drug-like.[2019optimization] Thus, a comprehensive consideration of the whole ligand structure and its properties is required to modulate multiple properties simultaneously.
@@
Here, the goal is to provide modified molecular structures to meet a given multi-property condition. We tested our model for three scenarios that practitioners might encounter in a lead optimization process: (1) increasing or decreasing ligand lipophilicity while maintaining its molecular size, (2) increasing drug-likeness while maintaining its molecular weight, and (3) alleviating the synthetic complexity of a ligand with fair bioactivities. The second and third scenarios are more challenging than the first, which requires more comprehensive consideration of the relationship between the properties of interest. Throughout Section 2.2, the following criteria are commonly employed to select test molecules: (1) a single neutral molecule, (2) molecular weight that falls within the range of 100 to 500, and (3) inclusion of at least one BRICS bond. In addition, to examine our model’s capability to find novel bioisosterism, we only used fragments from the test dataset for insertion, which were never adopted either as ground truths or negative samples for the insertion step in the training process.
@@
The overall property control results are exhibited in Table 1. For scenario 1, we selected 1,000 unique molecules whose log values are within the range of 1 to 5, regarded as moderate lipophilicity. Whether one should increase a molecule’s log or not depends on its specific application; for example, the biological target under consideration or the drug delivery plan are critical factors in making such a decision. Thus, we tested our model for both cases with two property control conditions: increasing log by 1 and decreasing log by 1 while maintaining overall molecule size in terms of molecular weight (MW). We note that log is a logarithm of the octanol-water partition coefficient, so the difference by 1 corresponds to a 10 times larger or smaller solubility ratio.[logp] We generated ten unique molecular structures for each original molecule according to the two property control conditions. For a few test molecules, DeepBioisostere generated fewer than ten molecular structures for each one since the BRICS-allowed insertion fragments were fewer than ten. The results are summarized in the first row on Table 1. We observed that the mean values of the molecular weight differences between an original molecule and its modified ones were almost negligible: +0.04 and -0.28, respectively. These values suggest that the molecular weights of most modified structures are different from the original ones by less than one heavy atom. (See Supplementary Figure 4 for more details about property distributions) Meanwhile, log was modulated with averaged changes of +0.83 and -0.84, falling slightly behind the requested property values. These statistics indicate that our model successfully adjusted the lipophilicity of the given molecular structures with barely changing their molecular weights, which requires careful selection of both removal and insertion fragments.
@@
| Scenario 1 | MW | 0 | +0.04 | 12.8 |
| log | +1 | +0.83 | 0.38 | |
| MW | 0 | -0.28 | 12.6 | |
| log | -1 | -0.84 | 0.43 | |
| Scenario 2 | MW | 0 | -0.11 | 12.1 |
| QED | +0.1 | +0.067 | 0.090 | |
| MW | 0 | -0.57 | 12.0 | |
| QED | +0.2 | +0.100 | 0.099 | |
| Scenario 3 | QED | 0 | -0.026 | 0.073 |
| SAscore | -0.5 | -0.424 | 0.433 | |
| QED | 0 | -0.028 | 0.074 | |
| SAscore | -1.0 | -0.719 | 0.524 | |
Second, we demonstrated an optimization process where a molecular structure is modified to improve its drug-likeness in terms of quantitative estimation of drug-likeness (QED).[qed] (Scenario 2 in Table 1) In this scenario, only molecules whose QED values are less than 0.4 were used as test molecules. Property control conditions were set to increase the QED value by 0.1 or 0.2 while keeping the molecular weight constant. By employing the DeepBioisostere model, we were able to obtain novel molecules with enhanced drug-likeness; the mean values of QED changes were +0.067 and +0.100, respectively. This result indicates that our model can modulate drug-likeness without much change in the original molecular structure, which is critical in order not to lose pre-established crucial properties. Although the property control condition specifying a larger QED increase yielded molecules with higher QED values, the deviation from the given condition also became larger. It is not surprising that increasing QED by 0.2 is even more challenging, as the occurrence of such replacements within our training dataset is exceedingly rare. Refer to supplementary Section 4.1 for further discussion about the statistics.
Lastly, we tested our model to enhance the synthetic accessibility of a molecule with a high QED value but low synthetic accessibility. (scenario 3 in Table 1) Here, we adopted the SAscore introduced by sascore, a widely used estimator of synthetic accessibility. We collected test molecules for this scenario with the criteria that the QED value is greater than 0.7 and the SAscore is greater than 4. Note that the SAscore of a simple catalog molecule is about 2 or 3, whereas for a natural product, it is mostly greater than 5.[sascore] Property control conditions were set to decrease the SAscores of molecules by 0.5 or 1.0 while retaining their QED values. In this scenario, our model provided novel molecules with lower SAscores, although several of them exhibited somewhat decreased QED values. This result suggests that DeepBioisostere can be adopted to supplement many in silico methods for drug design, whose outcomes are often questioned regarding their synthetic accessibility.[walters2018virtual, gao2020synthesizability]